# MiniTac: An Ultra-Compact Vision-Based Tactile Sensor for Enhanced Palpation in Robot-Assisted Minimally Invasive Surgery

### v1

@@ Abstract
In addressing this limitation, there has been significant interest in integrating tactile sensors into RAMIS systems [7, 8]. However, the trocar port size—typically an 8 mm diameter circle, as is standard in the widely-used Da Vinci robotic surgical systems [9]—presents a unique challenge. This challenge involves developing a sensor that not only provides high-density tactile sensing but is also compact enough for insertion through the port. Despite extensive research, current tactile sensing solutions continue to struggle with achieving miniaturization while maintaining a sufficient number of taxels (sensing units) within the limited space (reviewed in detail in Sec. II). This challenge has hindered the integration of tactile sensors into RAMIS systems and limited their overall impact. The lack of a commercially available solution further highlights the complexity of this challenge.
In this paper, we introduce MiniTac, a novel vision-based tactile sensor specifically designed for RAMIS systems. The ultra-compact sensor has a diameter of (Fig. 1(a)), ideally suited to fit through standard trocar ports of Da Vinci robotic surgical systems (Fig. 1(b)). MiniTac features a mechanoresponsive photonic elastomer membrane that alters its color distribution in response to different levels of contact [10]. This altered color distribution is captured by a miniature camera at a high spatial resolution of , endowing the device with sensitivity akin to having taxels across its surface. This enables MiniTac to accurately map detailed deformations at the contact site and reliably identify simulated embedded tumors within the contact region, as depicted in Fig. 1(c).
@@
The remainder of the paper is organized as follows: Sec. II reviews relevant literature. Sec. III introduces MiniTac, detailing its design objectives, fabrication processes, calibration methods, and sensitivity. Sec. IV validates the capabilities of MiniTac in detecting hard embedded tumors in both phantom models and ex-vivo tissue. In Sec. V, we discuss the unique features and advantages of MiniTac, emphasizing its potential to enhance palpation-equivalent capabilities in RAMIS as compared to other tactile sensors. We conclude with key findings and implications in Sec. VI.
@@ II Related Work
Vision-based tactile sensors utilize an embedded camera to capture optical changes of a front soft membrane, decoding its deformation and providing tactile feedback. These sensors offer extremely high spatial resolution at the pixel level, straightforward manufacturing processes, and low costs. Based on their working principles, they can be broadly categorized into photometric stereo types [16, 17, 18, 19, 20], darkness mapping types [21], binocular stereo types [22], and compound-eye stereo types [23]. Extensive research has demonstrated that these sensors can precisely detect object texture, hardness, and surface roughness, closely mirroring the efficacy of manual palpation [24, 25, 26]. Indeed, their effectiveness in identifying simulated embedded tumors has been demonstrated, highlighting their potential in medical applications [27, 28, 29]. However, these sensors tend to be bulky within the context of RAMIS, posing significant challenges for their integration into RAMIS systems. The characteristics of these sensors are summarized in Tab. I.
@@ III The MiniTac Sensor
This section outlines the developmental process for MiniTac, starting with design specifications that would ensure compatibility with RAMIS, as detailed in Sec. III-A. We then describe the design and fabrication processes developed to meet these specifications—see Sec. III-B. This leads to a discussion about the calibration method employed to correlate raw sensor readings with deformation data at the contact site, as outlined in Sec. III-C. Finally, we test MiniTac’s sensitivity, as detailed in Sec. III-D, and its repeatability and hysteresis in Sec. III-E.
@@ III-B Design and Fabrication Details
We carefully design and manufacture the MiniTac sensor to meet the given objectives and specifications, as illustrated in the exploded view and schematic diagram—see Fig. 2(a) and Fig. 2(b). Consideration is given to the design and selection of MiniTac’s components, as each element will have an impact on the overarching goal.
@@
Taking into account the stringent requirements in terms of physical dimensions and sensitivity to deformation, we opt for a mechanoresponsive photonic membrane, as illustrated in Miller et al. [10]. Unlike conventional color-changeable membranes [16], which necessitate complex and bulky lighting systems, this membrane boasts a remarkable feature: on deformation, it is susceptible to changes in reflection wavelength, leading to color change even under basic illumination. Images of the undeformed and deformed membrane are depicted in Fig. 3(a) and (b), respectively. We refer the readers to the original paper [10] for details of the principles that underlie the workings of this mechanoresponsive photonic membrane.
@@
where is an offset, typically set to , to convert negative differences to positive values, and is a clipping function ensuring the value falls within the image range , so as to be compatible with the RGB value range for visualization. The augmented tactile imprint of the contact shown in Fig. 3(b) can be seen, more clearly, in Fig. 3(c).
@@
Because the membrane’s reflective surface can cause significant glare, (see illustration in Fig. 3(d)), we incorporate a diffusing layer made of translucent silicone, specifically Posilicone DRSGJ02, positioned directly behind the membrane. The chosen material not only surpasses traditional methods such as diffuser paper in terms of light distribution uniformity but also adheres naturally to the photonic membrane. This adhesive property is beneficial as it mitigates the use of additional bonding agents whose constituent chemicals could interact with the membrane causing color alterations or other unwanted issues. The translucent quality of the silicone ensures even light distribution, enhancing the sensor’s accuracy in detecting and responding to color changes.
@@
To enhance the sensor’s robustness and durability, we designed a metal shell with an outer diameter of and a minimal wall thickness of , targeting a large sensing region during RAMIS. This shell securely houses the camera module as shown in Fig. 2. The lower end of the shell is attached to a carbon tube via a custom fixture, which acts as an extended handle. The design strengthens the sensor’s structure while enhancing its usability and ergonomic handling, ultimately facilitating its seamless integration into RAMIS systems.
@@ III-C Calibration
We analyze and illustrate these correlations to understand the relationship between membrane deformation and the HSV channels, as shown in Fig. 4. Ideally, a uniform membrane color would enable straightforward mapping from color variation to surface deformation using a single function. However, real-world variations in membrane color, as observed in Fig. 3, necessitate a more complex approach. To address this, we introduce positional factors into our calibration model. We define a function that maps color changes at each pixel position to the corresponding deformation depth. The function is mathematically formulated as follows:
@@
Fig. 5 presents qualitative results, illustrating the reconstructed sensor surface deformation from various kinds of contact, demonstrating our calibration method’s efficacy.
@@ III-D Sensor Sensitivity
To evaluate the sensitivity of MiniTac, we use an ATI Nano17 force sensor to apply controlled pressure to the sensor’s surface and record its readings (Fig. 6(a)). As shown in Fig. 6(b), the sensor demonstrates high sensitivity, is capable of detecting deformations at , can resolve force increments at approximately , and does not saturate under forces up to .
@@ III-E Sensor Repeatability and Hysteresis
With the same setting in Fig. 6(a), the repeatability of MiniTac is calculated as:
@@
The red curve in Fig. 7 is the smoothed average of the five trials measured above. Based on this curve, the maximum measurement difference between two processes is observed as , and the hysteresis is calculated as
@@ IV Experiments
Initially, we evaluate the efficacy of MiniTac in detecting hard embedded tumors using phantoms (Sec. IV-A), establishing a decision boundary so as to distinguish tumor tissue from normal tissue. Further validation is then performed ex-vivo to verify the efficacy of MiniTac as well as the established decision boundary (Sec. IV-B).
@@ IV-A Experiments on Phantoms
To create phantoms that closely resemble human tissue, we use Smooth-On Ecoflex 00-30 silicone rubber, which has a Shore 00 hardness of 30, similar to that of human tissue. The methodology follows the guidelines described in [27]. To assess the performance of MiniTac in a variety of tumor scenarios, we embed metal balls of assorted diameters across different depths , simulating a range of embedded tumor situations. The phantom setup is illustrated in Fig. 8.
@@
To precisely control the force applied during data collection, we construct a specialized device as depicted in Fig. 9. It features a platform that securely holds the MiniTac in place and also allows for the placement of weights to regulate the applied force. The movement of the platform is constrained by two smooth vertical shafts, ensuring that the force is directed strictly downwards.
One example for each scenario is showcased in Fig. 10(b) for positive samples and Fig. 10(c) for negative samples, respectively. The data from positive samples (Fig. 10(b)) distinctly indicate the presence of embedded tumors, with larger diameters and shallower depths resulting in more pronounced deformation signatures. In contrast, Fig. 10(c) demonstrates that deformations in phantoms without tumors tend to have flatter features due to the uniformity of the phantom material.
@@ IV-B Experiments on Ex-Vivo Tissues
Fig. 11 illustrates our ex-vivo testing process. In Fig. 11(a), we show an ex-vivo tissue sample with areas containing and lacking an embedded tumor, as identified by experienced surgeons through manual palpation. Using MiniTac, surgeons then performed 50 random presses across this tissue. Our established decision boundary (Eq. 5) achieved 100% accuracy in these trials, perfectly aligning with the surgeons’ ground truth assessments. Fig. 11(b) provides two representative samples of reconstructed deformation and classification results. This strong correlation between MiniTac’s outputs and expert surgical evaluations demonstrates the sensor’s effectiveness in accurately detecting tumors within human tissue.
@@ V Discussion
The ability to accurately map deformations and calculate mean and variance has proven effective in the detection of hard objects such as tumors, as demonstrated in controlled phantom experiments (Sec. IV-A) and ex-vivo tissue applications (Sec. IV-B). Harder objects induce more pronounced deformations, enhancing detectability and increasing the deformation data’s mean and variance. The decision boundary analysis confirms the importance of standard deviation, which remains significant even when larger forces are applied (Eq. 5). Although similar results might be achieved with high-resolution vision-based tactile sensors such as GelSight[27] and DIGIT[29], MiniTac’s simple illumination system allows for a significantly more compact design, facilitating integration into RAMIS systems (Fig. 1).

### v2

@@ Abstract
In addressing this limitation, there has been significant interest in integrating tactile sensors into RAMIS systems [7, 8]. However, the trocar port size—typically an 8 mm diameter circle, as is standard in the widely-used Da Vinci robotic surgical systems [9]—presents a unique challenge. This challenge involves developing a sensor that not only provides high-density tactile sensing but is also compact enough for insertion through the port. Despite extensive research, current tactile sensing solutions continue to struggle with achieving miniaturization while maintaining a sufficient number of taxels (sensing units) within the limited space (reviewed in detail in Sec.˜II). This challenge has hindered the integration of tactile sensors into RAMIS systems and limited their overall impact. The lack of a commercially available solution further highlights the complexity of this challenge.
In this paper, we introduce MiniTac, a novel vision-based tactile sensor specifically designed for RAMIS systems. The ultra-compact sensor has a diameter of (Fig.˜1(a)), ideally suited to fit through standard trocar ports of Da Vinci robotic surgical systems (Fig.˜1(b)). MiniTac features a mechanoresponsive photonic elastomer membrane that alters its color distribution in response to different levels of contact [10]. This altered color distribution is captured by a miniature camera at a high spatial resolution of , endowing the device with sensitivity akin to having taxels across its surface. This enables MiniTac to accurately map detailed deformations at the contact site and reliably identify simulated embedded tumors within the contact region, as depicted in Fig.˜1(c).
@@
The remainder of the paper is organized as follows: Sec.˜II reviews relevant literature. Sec.˜III introduces MiniTac, detailing its design objectives, fabrication processes, calibration methods, and sensitivity. Sec.˜IV validates the capabilities of MiniTac in detecting hard embedded tumors in both phantom models and ex-vivo tissue. In Sec.˜V, we discuss the unique features and advantages of MiniTac, emphasizing its potential to enhance palpation-equivalent capabilities in RAMIS as compared to other tactile sensors. We conclude with key findings and implications in Sec.˜VI.
@@ II Related Work
Vision-based tactile sensors utilize an embedded camera to capture optical changes of a front soft membrane, decoding its deformation and providing tactile feedback. These sensors offer extremely high spatial resolution at the pixel level, straightforward manufacturing processes, and low costs. Based on their working principles, they can be broadly categorized into photometric stereo types [16, 17, 18, 19, 20], darkness mapping types [21], binocular stereo types [22], and compound-eye stereo types [23]. Extensive research has demonstrated that these sensors can precisely detect object texture, hardness, and surface roughness, closely mirroring the efficacy of manual palpation [24, 25, 26]. Indeed, their effectiveness in identifying simulated embedded tumors has been demonstrated, highlighting their potential in medical applications [27, 28, 29]. However, these sensors tend to be bulky within the context of RAMIS, posing significant challenges for their integration into RAMIS systems. The characteristics of these sensors are summarized in Tab.˜I.
@@ III The MiniTac Sensor
This section outlines the developmental process for MiniTac, starting with design specifications that would ensure compatibility with RAMIS, as detailed in Sec.˜III-A. We then describe the design and fabrication processes developed to meet these specifications—see Sec.˜III-B. This leads to a discussion about the calibration method employed to correlate raw sensor readings with deformation data at the contact site, as outlined in Sec.˜III-C. Finally, we test MiniTac’s sensitivity, as detailed in Sec.˜III-D, and its repeatability and hysteresis in Sec.˜III-E.
@@ III-B Design and Fabrication Details
We carefully design and manufacture the MiniTac sensor to meet the given objectives and specifications, as illustrated in the exploded view and schematic diagram—see Fig.˜2(a) and Fig.˜2(b). Consideration is given to the design and selection of MiniTac’s components, as each element will have an impact on the overarching goal.
@@
Taking into account the stringent requirements in terms of physical dimensions and sensitivity to deformation, we opt for a mechanoresponsive photonic membrane, as illustrated in Miller et al. [10]. Unlike conventional color-changeable membranes [16], which necessitate complex and bulky lighting systems, this membrane boasts a remarkable feature: on deformation, it is susceptible to changes in reflection wavelength, leading to color change even under basic illumination. Images of the undeformed and deformed membrane are depicted in Fig.˜3(a) and (b), respectively. We refer the readers to the original paper [10] for details of the principles that underlie the workings of this mechanoresponsive photonic membrane.
@@
where is an offset, typically set to , to convert negative differences to positive values, and is a clipping function ensuring the value falls within the image range , so as to be compatible with the RGB value range for visualization. The augmented tactile imprint of the contact shown in Fig.˜3(b) can be seen, more clearly, in Fig.˜3(c).
@@
Because the membrane’s reflective surface can cause significant glare, (see illustration in Fig.˜3(d)), we incorporate a diffusing layer made of translucent silicone, specifically Posilicone DRSGJ02, positioned directly behind the membrane. The chosen material not only surpasses traditional methods such as diffuser paper in terms of light distribution uniformity but also adheres naturally to the photonic membrane. This adhesive property is beneficial as it mitigates the use of additional bonding agents whose constituent chemicals could interact with the membrane causing color alterations or other unwanted issues. The translucent quality of the silicone ensures even light distribution, enhancing the sensor’s accuracy in detecting and responding to color changes.
@@
To enhance the sensor’s robustness and durability, we designed a metal shell with an outer diameter of and a minimal wall thickness of , targeting a large sensing region during RAMIS. This shell securely houses the camera module as shown in Fig.˜2. The lower end of the shell is attached to a carbon tube via a custom fixture, which acts as an extended handle. The design strengthens the sensor’s structure while enhancing its usability and ergonomic handling, ultimately facilitating its seamless integration into RAMIS systems.
@@ III-C Calibration
We analyze and illustrate these correlations to understand the relationship between membrane deformation and the HSV channels, as shown in Fig.˜4. Ideally, a uniform membrane color would enable straightforward mapping from color variation to surface deformation using a single function. However, real-world variations in membrane color, as observed in Fig.˜3, necessitate a more complex approach. To address this, we introduce positional factors into our calibration model. We define a function that maps color changes at each pixel position to the corresponding deformation depth. The function is mathematically formulated as follows:
@@
Fig.˜5 presents qualitative results, illustrating the reconstructed sensor surface deformation from various kinds of contact, demonstrating our calibration method’s efficacy.
@@ III-D Sensor Sensitivity
To evaluate the sensitivity of MiniTac, we use an ATI Nano17 force sensor to apply controlled pressure to the sensor’s surface and record its readings (Fig.˜6(a)). As shown in Fig.˜6(b), the sensor demonstrates high sensitivity, is capable of detecting deformations at , can resolve force increments at approximately , and does not saturate under forces up to .
@@ III-E Sensor Repeatability and Hysteresis
With the same setting in Fig.˜6(a), the repeatability of MiniTac is calculated as:
@@
The red curve in Fig.˜7 is the smoothed average of the five trials measured above. Based on this curve, the maximum measurement difference between two processes is observed as , and the hysteresis is calculated as
@@ IV Experiments
Initially, we evaluate the efficacy of MiniTac in detecting hard embedded tumors using phantoms (Sec.˜IV-A), establishing a decision boundary so as to distinguish tumor tissue from normal tissue. Further validation is then performed ex-vivo to verify the efficacy of MiniTac as well as the established decision boundary (Sec.˜IV-B).
@@ IV-A Experiments on Phantoms
To create phantoms that closely resemble human tissue, we use Smooth-On Ecoflex 00-30 silicone rubber, which has a Shore 00 hardness of 30, similar to that of human tissue. The methodology follows the guidelines described in [27]. To assess the performance of MiniTac in a variety of tumor scenarios, we embed metal balls of assorted diameters across different depths , simulating a range of embedded tumor situations. The phantom setup is illustrated in Fig.˜8.
@@
To precisely control the force applied during data collection, we construct a specialized device as depicted in Fig.˜9. It features a platform that securely holds the MiniTac in place and also allows for the placement of weights to regulate the applied force. The movement of the platform is constrained by two smooth vertical shafts, ensuring that the force is directed strictly downwards.
One example for each scenario is showcased in Fig.˜10(b) for positive samples and Fig.˜10(c) for negative samples, respectively. The data from positive samples (Fig.˜10(b)) distinctly indicate the presence of embedded tumors, with larger diameters and shallower depths resulting in more pronounced deformation signatures. In contrast, Fig.˜10(c) demonstrates that deformations in phantoms without tumors tend to have flatter features due to the uniformity of the phantom material.
@@ IV-B Experiments on Ex-Vivo Tissues
Fig.˜11 illustrates our ex-vivo testing process. In Fig.˜11(a), we show an ex-vivo tissue sample with areas containing and lacking an embedded tumor, as identified by experienced surgeons through manual palpation. Using MiniTac, surgeons then performed 50 random presses across this tissue. Our established decision boundary (Eq.˜5) achieved 100% accuracy in these trials, perfectly aligning with the surgeons’ ground truth assessments. Fig.˜11(b) provides two representative samples of reconstructed deformation and classification results. This strong correlation between MiniTac’s outputs and expert surgical evaluations demonstrates the sensor’s effectiveness in accurately detecting tumors within human tissue.
@@ V Discussion
The ability to accurately map deformations and calculate mean and variance has proven effective in the detection of hard objects such as tumors, as demonstrated in controlled phantom experiments (Sec.˜IV-A) and ex-vivo tissue applications (Sec.˜IV-B). Harder objects induce more pronounced deformations, enhancing detectability and increasing the deformation data’s mean and variance. The decision boundary analysis confirms the importance of standard deviation, which remains significant even when larger forces are applied (Eq.˜5). Although similar results might be achieved with high-resolution vision-based tactile sensors such as GelSight[27] and DIGIT[29], MiniTac’s simple illumination system allows for a significantly more compact design, facilitating integration into RAMIS systems (Fig.˜1).